# Using matrix-product states for time-series machine learning

### v1

@@ IV MPS interpretability
Figure 7: Sampling MPS joint probability distributions. At each time-point, the time-marginalized kernel density estimate (KDE) of time-series amplitudes is visualized as a heat map, overlaid with individual training instances for reference. For each dataset – (a) ECG, (b) Power Demand, and (c) the most complex Noisy, Trendy Sinusoid (NTS) dataset from Sec. III.2.1 with 3 periods and 3 trends – we show two panels: in subplot (i), we show the ground-truth training data, with the KDE heat map computed from the training instances and overlaid with examples of time series from the training set. Subplot (ii) shows the synthetic data generated by sampling from a $\chi_{\rm max}=80$, $d=12$ MPS trained on the corresponding dataset. Using inverse transform sampling, we generate each time-series amplitude conditionally from the MPS-encoded distribution, computing the KDE heat map from the synthetic dataset, and overlaying with representative examples of the synthetic single-shot trajectories.
A compelling justification for using MPS to infer complex joint time-series distributions directly from data lies in their inherent interpretability, which sets them apart from ‘black-box’ approaches. For example, the single-site reduced density matrix (RDM) $\rho$, as in Eq. 14, can be used to extract the learned conditional probability distribution of time-series amplitudes at a given time point $t$, directly from the MPS. If amplitudes of other sites are known, this can be used to constrain the distribution given by $\rho$ by conditioning other sites on their known values. As we have shown in Sec. III.2, the RDM forms the foundation of conditional site-wise sampling, enabling us to effectively impute missing values in time series by leveraging the temporal dependencies encoded by the MPS during training. Beyond its role in sampling, the RDM, which encapsulates all interactions between a given subsystem and the rest of the MPS, serves as a powerful tool for characterizing the local and global correlation structures captured within the model.
@@
In this section, we harness analytical tools and machinery from quantum information theory to showcase the rich interpretability of MPSTime. Specifically, we demonstrate how these tools provide meaningful insights into the complex temporal correlations learned by the MPS. In Sec. IV.1, we study the generative distribution learned by the MPS through sampling entire time-series trajectories from the encoded joint distribution and comparing these to real-world instances from the training set. Through this qualitative investigation, we aim to determine whether the MPS is able to successfully capture complex temporal correlations from a finite number of training instances, and as a result, produce plausible synthetic data, as the sampling process explicitly relies on the temporal dependencies learned at each MPS site. In Sec. IV.2, we show how the complex correlations encoded by the MPS can be understood through the lens of entanglement entropy, extending fundamental concepts from quantum information theory – traditionally applied to analyze MPS in quantum many-body physics – to the analysis of temporal correlation structures learned by the MPS in a time-series ML setting.
IV.1 Unveiling temporal correlation structures learned by MPS through sampling
Building upon our earlier results in Sec. III.2, where we used partially observed time series to impute missing values from the conditional distribution represented by the MPS, we now study how accurately MPSTime can capture the joint time-series distribution from which a given time-series dataset is sampled. To this end, we generated single-shot trajectories (entire $T$-sample instances) directly from the MPS-encoded joint distribution. Unlike imputation, which begins with conditioning the MPS on partially observed time series, we randomly sample values directly from the unconditioned joint distribution using inverse transform sampling. The sampling process begins at the first MPS site, then proceeds sequentially from left-to-right, at each site, selecting a random state from the RDM, then updating the MPS to reflect the selected state, until all sites have been sampled once. This yields a single time-series instance of $T$ samples. By returning to the first site and sampling again, additional independent trajectories can be generated from the modeled joint distribution. A discussion of the specific inverse transform sampling method we use is given in Appendix E.
@@ IV.1 Unveiling temporal correlation structures learned by MPS through sampling
For our investigation, we focused on one synthetic (NTS6) and two empirical datasets (Power Demand and ECG). For the purposes of qualitative comparison, we selected values of $\chi_{\rm max}$ and $d$ that ensured the MPS possessed sufficient expressibility to model the essential temporal correlations (i.e., larger $\chi_{\rm max}$), while producing trajectories from the encoded generative distribution with small encoding error (i.e., larger $d$). To achieve this, we trained a $\chi_{\rm max}=85$ and $d=12$ MPS for all three datasets. From each MPS, we then generated the same number of trajectories as there were time-series instances in the training set, allowing for a direct comparison between realizations from the various modeled and original data distributions. In Fig. 7 we qualitatively compare the synthetically generated time-series instances from the MPS to training instances from the ground-truth joint distribution. In all cases, the MPS was able to replicate the data distribution underlying a finite number of time-series instances, as evidenced by its ability to generate new time-series that closely resemble the training data.
@@
Figure 8: Interpreting the joint probability distribution of the MPS using the single-site entanglement entropy (SEE). By sequentially updating the MPS ($d=12$, $\chi_{\rm max}=80$) with each additional measured site, the SEE reveals how the entanglement in the remaining unmeasured sites evolves. This SEE structure provides insight into the contextual dependencies captured by the MPS, which, as it is conditioned on each additional site, reflects an evolving conditional distribution. Panel (a) shows the mean SEE across the Power Demand test set as a function of the number of sequentially measured sites (y-axis), starting from the first site and progressing from left (site 1) to right (site 24). The accompanying bar plot depicts the residual entanglement entropy – defined as the average unmeasured SEE – for varying numbers of measured sites. Panel (b) shows representative time-series trajectories sampled from the MPS for varying numbers of measured sites. These trajectories demonstrate how the conditional distribution of time-series amplitudes evolves as increasingly many points are measured: (i) no sites measured, (ii) 4 sites measured, (iii) 12 sites measured, (iv) 18 sites measured. The solid red line indicates the ground-truth time-series being used to condition the MPS.
IV.2 Conditional entanglement entropy
@@
In a quantum mechanical system, the single-site entanglement entropy (SEE) quantifies the amount of entanglement shared between a single site and the remainder of the system [28]. When adapted to MPS-based time-series ML, the SEE determines the extent to which the time-series amplitude at a single site depends on the value of all other amplitudes (see Fig. 1(c)). If the SEE is large, this implies strong correlations with the remainder of the MPS, i.e., determining the amplitudes at other sites (time points) can reduce uncertainty about the amplitude at the site under consideration. On the other hand, a small SEE suggests the site behaves more independently, and therefore, determining the amplitudes of other sites does little to constrain its value through uncertainty reduction.
@@
Examining how the SEE changes after conditioning on the time-series amplitude at one or more sites can therefore provide insight into the conditional relationships learned by the MPS. As a concrete example, consider an unmeasured site $B$ with $\textrm{SEE}_{\rm before}>0$. We then “measure” another site $A$, projecting the MPS onto a subspace in which the local state at $A$ is fixed to the observed time-series value. If this conditioning reduces the SEE at site $B$, i.e., $\textrm{SEE}_{\rm after}<\textrm{SEE}_{\rm before}$, this implies that $B$’s state is strongly influenced by knowledge of $A$’s state. In other words, observing a particular time-series amplitude $x_{A}$ at site $A$ constrains the distribution of amplitudes $x_{B}$ can take at site $B$. Extending this approach across multiple pairs or sets of sites can allow us to ‘map out’ more complex correlation structures encoded by the MPS.
@@
where $\lambda_{k}$ are the eigenvalues of the single-site reduced density matrix $\rho_{i}$ with rank $r$.
@@
To showcase the insights facilitated by the SEE, we analyzed the same MPS trained on the Power Demand dataset in Sec. IV.1. For our analysis, we aimed to uncover the conditional dependencies learned by the MPS by observing how its entanglement structure evolves as increasingly many site-wise measurements are made. Here, we chose to perform measurements in a sequential order, starting at the first site and incrementing the number of measured sites by one. With each additional measurement, we analyzed the SEE of the remaining unmeasured sites in order to isolate the influence of the measurement on the resulting entanglement structure.
@@
In Fig. 8(a)(i), we visualize the conditional entanglement structure of the MPS by averaging the measurement-dependent SEE over the entire Power Demand test set of $1029$ instances. As a proxy for the remaining entanglement encapsulated by the MPS after cumulative measurements, we also computed the residual entanglement entropy as the average SEE of the unmeasured sites. As expected, as more sites of the MPS are conditioned, the SEE decreases throughout the MPS. Interestingly, there appear to be key sites where, after they are conditioned, the SEE drops sharply. This occurs after measuring sites (correspond to time values in the MPS) 1, 2, 8, and 18. This indicates that the values at these points are likely crucial in determining the values of other points within the MPS. Figure 8(b) shows the sampled trajectories of the joint probability distribution learned by the MPS conditioned on 0-18 sites being measured with known amplitudes of a single time series. Moving from (i)-(iv) we clearly observe a reduction in the fluctuations in the trajectory amplitudes as more sites are measured. The reduction in fluctuations is particularly prominent between $t=15$ and $t=23$. This indicates that the conditional SEE can be used as a proxy for determining the certainty of the values of unmeasured sites conditioned on the measurement outcomes of other sites. The results illustrated in Fig. 8 highlight the utility of the MPSTime, which can be used to compute the conditional trajectories and conditional SEE. We emphasize that such calculations would be highly challenging or impossible to perform with the empirical dataset alone, while they are naturally extracted from the MPS framework.

### v2

@@ IV MPS interpretability
Figure 7: Sampling MPS joint probability distributions over time series yields synthetic examples that closely resemble the training data. At each time-point, the time-marginalized kernel density estimate (KDE) of time-series amplitudes is visualized as a heat map, and is overlaid with individual time-series instances from the entire training set (white traces). The color intensity within each heatmap indicates the time-series amplitude density relative to the maximum at each time-point. For each dataset – (a) ECG, (b) Power Demand, and (c) the most complex Noisy, Trendy Sinusoid dataset (NTS5) from Sec. III.2.1 with 3 periods and 3 trends – we show an upper and a lower panel: in subplot (i), we show the ground-truth training data as white traces, with the KDE computed from the full training set across all data classes. Subplot (ii) shows the synthetic data generated by sampling from a sufficiently expressive $\chi_{\rm max}=80$, $d=18$ MPS trained on the same dataset. Using inverse transform sampling (see Appendix E), we generate each time-series amplitude conditionally from the MPS-encoded distribution, and depict representative examples of the resulting trajectories with white traces.
A compelling justification for using matrix-product states to infer joint distributions directly from time-series data lies in their inherent interpretability, which sets them apart from ‘black-box’ approaches. By inferring a closed-form approximation of the joint distribution, MPSTime permits analysis of the underlying time-series distribution that would be computationally intractable for standard methods. In this section, we showcase two aspects of interpretable MPS-based machine learning that are enabled by studying the single-site reduced density matrix (RDM) $\rho$ in Eq. (14). At its core, the single site RDM captures both the local statistical properties of the time-series at a given point. Crucially, the RDM is easily extractable directly from the MPS, and contains all the information needed to sample from the approximated joint distribution, as well as analyze the short-range and long-range temporal correlation structures that it encodes.
@@
First, in Sec. IV.1, by comparing samples generated by the MPS to the training dataset, we show that MPSTime can successfully capture complex temporal correlations from a finite number of training instances. This indicates that MPSTime can be used to generate new synthetic time-series examples from any given data class, an application that has broad implications [70, 71, 72]. Second, in Sec. IV.2, we show how the complex correlations encoded by the MPS can be understood through the lens of entanglement entropy, extending fundamental concepts from quantum information theory – traditionally applied to analyze MPS in quantum many-body physics – to the analysis of temporal correlation structures learned by MPSTime. Third, in Sec. IV.3, we study the ability of MPSTime to generalize learned data structures to new, out-of-distribution (OOD) contexts.
IV.1 Qualitatively evaluating the learned joint distribution
In Sec. III.2, we demonstrated that MPSTime accurately imputes missing values in partially observed time series by sampling from its learned joint distribution, conditioned on observed values. Here, we extend this analysis to examine the full joint distribution encoded by the MPS, providing insight into what temporal correlations and patterns MPSTime learns from the training data. Instead of conditioning on observed values, as in Sec. III.2, we now generate entire time-series instances by sampling values directly from the unconditioned MPS. For clarity, we refer to MPS-generated data as ‘trajectories’, and real training data as ‘time-series instances’. Our sampling procedure follows the natural temporal ordering of time series, beginning at the first MPS site ($t=1$) and advancing sequentially until reaching the final site ($t=T$). Each time, a random state is selected from the reduced density matrix (RDM) using inverse transform sampling, and the MPS is updated according to the sampled value. This site-wise process of sampling and updating the MPS continues until generating a complete trajectory of $T$ samples. By returning to $t=1$ and sampling again, additional trajectories can be generated from the MPSTime-modeled joint distribution. For implementation details of our inverse transform sampling approach, which is similar to recent work by Mossi et al. [21], see Appendix E.
@@ IV.1 Qualitatively evaluating the learned joint distribution
For our investigation, we focused on one synthetic dataset (NTS5) and two empirical datasets (Power Demand and ECG) from Sec. III.2. As with imputation, we discarded class-specific information for each dataset, where applicable, by treating each time-series instance as belonging to the same data class. For the purposes of qualitative comparison, we selected values of $\chi_{\rm max}$ and $d$ that ensured the MPS was sufficiently expressive to model the essential temporal correlations (i.e., larger $\chi_{\rm max}$), while producing trajectories from the encoded generative distribution with small encoding error (i.e., larger $d$). To achieve this, we trained a $\chi_{\rm max}=80$, $d=18$ MPS for all three datasets. From each MPS, we then generated the same number of trajectories as there were time-series instances in the training set, allowing for a direct comparison between realizations from the various modeled and original data distributions.
In Fig. 7 we qualitatively compare the synthetically generated trajectories from MPSTime (lower panels) to training instances from the ground-truth joint distribution (upper panels) for each of the three datasets. In general, for each real and synthetic dataset pair, we observe that the lighter regions of the synthetic dataset KDEs (which indicate higher relative density) overlap closely with the typical amplitude ranges seen in the ground-truth datasets. This correspondence between the relative densities of the synthetic and real time-series datasets at each time point suggests that MPSTime has accurately approximated the underlying distribution. Inspecting the individual time series in the upper panels, shown as superimposed white traces, we observe that the MPS-generated trajectories are comparatively less smooth than those from the real (training) dataset, particularly for the longer ECG and NTS5 time series. This effect is to be expected, as the encoding and approximation errors inherent in the MPS representation, as discussed in Appendix B, limit the precision with which finer temporal features can be resolved.
Across all three datasets – ECG, Power Demand, and NTS5 – MPSTime captures essential qualitative temporal features present in the training data. For the ECG dataset in Fig. 7(a), synthetic trajectories in (a-ii) replicate the characteristic cardiac cycle of the training data in (a-i), including the pronounced QRS complex (initial peak) near $t=1$ and the more subtle secondary T wave near $t=80$.
For the Power Demand dataset in Fig. 7(b), the generated trajectories in Fig. 7(b-ii) capture the two-peak structure and intervening dip present in Fig. 7(b-i), mirroring the cyclic patterns of real power usage over the 24 hour period. Compared to the training data, the Power Demand trajectories are less diverse. Notably, very few trajectories begin with values above $x=0$ at $t=1$ or reach values exceeding $x=2$ around $t=22$. These discrepancies may arise from outlier values being under-represented by the MPS approximation of the joint distribution, an effect that is especially pronounced for small datasets such as IPD with only $67$ training instances. Further, the diversity of trajectories generated by our rejection-based sampling approach – implemented to avoid sampling unphysical states due to artificial distribution broadening (see Appendix. E) – further reduces the likelihood of observing these rare outlier values.
Finally, for the NTS5 dataset in Fig. 7(c), the training data in Fig. 7(c-i) exhibits a particularly complex temporal structure arising from sinusoids with multiple trends, periods and randomized phases, as described in Sec. III.1.1. By superimposing time-series, shown as white traces, the training distribution reveals three distinct bands, corresponding to the three possible slopes ($m=-3,0,3$ in Eq. (22)) of noisy trendy sinusoid. The MPS-generated trajectories in Fig. 7(c-ii) largely reproduce these qualitative features by capturing the multi-period oscillatory and trend behavior, as well as the general amplitude ranges. However, the synthetic dataset shows comparatively less distinct band separation and a more diffuse relative density in the KDE heatmap, particularly for $t<20$. Once again, this is likely due to the MPS approximation and encoding error, which results in noisier data that blur the boundaries between bands, reducing the resolution of finer temporal features, and leading to a more diffuse density structure than its real counterpart in Fig. 7(c-i).
Overall, our findings suggest that MPSTime effectively replicates the data distribution underlying a finite number of time-series by generating synthetic data that closely resemble training data in their essential qualitative features. This provides evidence that MPSTime learns meaningful temporal correlations and structure from the time-series data. It should be noted that while synthetic trajectories were generally less smooth than their real counterparts, this can be mitigated by increasing the physical dimension $d$ of the MPS (see Appendix B), although doing so would incur higher computational costs. While we limit the scope of our investigation to a qualitative comparison of temporal features, future work could quantify distributional similarity through low dimensional projections, classifier-based discrimination, or metrics suited to high-dimensional spaces, among others [73].
@@
Figure 8: Interpreting the joint probability distribution encoded by a trained MPS using the single-site entanglement entropy (SEE). By sequentially updating a trained MPS ($d=12$, $\chi_{\rm max}=80$) with each additional measured site, the SEE reveals how the entanglement in the remaining unmeasured sites evolves. This SEE structure provides insight into the contextual dependencies captured by the MPS, which, as it is conditioned on each additional site, reflects an evolving conditional distribution. Panel (a) shows the mean SEE, with magnitude indicated by the color intensity, across the Power Demand test set as a function of the number of sequentially measured sites (vertical axis), starting from the first site and progressing from left ($t=1$) to right ($t=24$). The accompanying bar plot depicts the residual entanglement entropy – defined as the average unmeasured SEE – for varying numbers of measured sites. Panel (b) shows representative time-series trajectories sampled from the MPS for varying numbers of measured sites. These trajectories demonstrate how the conditional distribution of time-series amplitudes evolves as increasingly many points are measured, for: (i) no sites measured; (ii) 4 sites measured; (iii) 12 sites measured; and (iv) 18 sites measured. The red curve in each panel depicts an unseen time-series instance from the test set for which time points lying to the left of the vertical dashed line (shaded window) have been used to condition the MPS.
IV.2 Extracting conditional entanglement entropy from the MPS
@@
In a quantum mechanical system, the single-site entanglement entropy (SEE) quantifies the amount of entanglement shared between a single site and the remainder of the system [28]. When adapted to an MPS trained on a time-series dataset, the SEE determines the extent to which the time-series amplitude at a single site depends on the value of all other amplitudes (see Fig. 1(c)). If the SEE is large, this implies strong correlations with the remainder of the MPS, i.e., determining the amplitudes at other sites (time points) can reduce uncertainty about the amplitude at the site under consideration. On the other hand, a small SEE suggests the site behaves more independently, and therefore, determining the amplitudes of other sites does little to constrain its value through uncertainty reduction.
@@
Examining how the SEE changes after conditioning on the time-series amplitude at one or more sites can therefore provide insight into the conditional relationships learned by the MPS. As a concrete example, consider an unmeasured site $B$ with $\textrm{SEE}_{\rm before}>0$. We then ‘measure’ another site $A$, projecting the MPS onto a subspace in which the local state at $A$ is fixed to the observed time-series value. If this conditioning reduces the SEE at site $B$, i.e., $\textrm{SEE}_{\rm after}<\textrm{SEE}_{\rm before}$, this implies that $B$’s state is strongly influenced by the local state at $A$. In other words, observing a particular time-series amplitude $x_{A}$ at site $A$ constrains the distribution of amplitudes $x_{B}$ can take at site $B$. Extending this approach across multiple pairs or sets of sites can allow us to ‘map out’ more complex correlation structures encoded by the MPS.
@@
where $\lambda_{k}$ are the eigenvalues of the single-site reduced density matrix $\rho_{i}$ with rank $r$. The single-site density matrix is always positive semi-definite and therefore the non-zero eigenvalues $\lambda_{k}>0$.
@@
To showcase the insights facilitated by the SEE, we analyzed the same MPS trained on the Power Demand dataset in Sec. IV.1. We aimed to uncover the conditional dependencies learned by the MPS by observing how its entanglement structure evolves as increasingly many site-wise measurements are made. We chose to perform measurements in a sequential order, starting at the first site and incrementing the number of measured sites by one. With each additional measurement, we analyzed the SEE of the remaining unmeasured sites in order to isolate the influence of the measurement on the resulting entanglement structure.
@@
In Fig. 8(a)(i), the conditional entanglement structure of the 24-site MPS is plotted as the average of the measurement-dependent SEE over the entire Power Demand test set (containing $1029$ time series). As a proxy for the remaining entanglement encapsulated by the MPS after cumulative measurements, we also computed the residual entanglement entropy as the average SEE of the remaining unmeasured sites. As expected, as more sites of the MPS are conditioned on known values, the SEE decreases. Interestingly, there appear to be key sites where, after they are conditioned, the SEE drops sharply. In particular, measuring sites (corresponding to time points) $t=1,2,8,18$, result in comparatively large reductions in SEE, indicating that the values at these points are crucial in determining the value of other points within the MPS. Figure 8(b) shows the sampled trajectories of the joint probability distribution learned by the MPS, after conditioning the MPS on the first 0 to 18 sites, using known values from a selected training instance. Moving from Figs 8(b)(i)–(iv), we observe a clear reduction in the fluctuations in the trajectory amplitudes as more sites are measured. This reduction is particularly prominent between $t=15$ and $t=23$. This indicates that the conditional SEE can be used as a proxy for estimating the uncertainty of unmeasured sites, conditioned on the measurement outcomes of other sites. The results illustrated in Fig. 8 highlight the utility of MPSTime, which can be used to compute the conditional trajectories and conditional SEE.
Figure 9: Evaluating the out-of-distribution generalization of MPSTime on noisy trendy sinusoids (NTS) with unobserved phases. MPSTime was trained on NTS (Eq.(22)) with eight discrete phases (indicated by the vertical dashed lines in the upper plot). To test generalization, we swept the phase parameter $\psi$ in Eq.(22) across the entire range $[0,2\pi)$ and imputed 50% missing data in unseen NTS with phase $\psi$. The Mean Absolute Error (MAE) between imputed and original noisy ground-truth values is shown for each value of $\psi$. The lower plot zooms into the range $[3\pi/4,\pi]$ to highlight the performance on intermediate phases. We compare the performance of MPSTime (red) with a 1-Nearest Neighbor imputation (1-NNI) baseline (gold), which cannot generalize beyond its training dataset. We also show a noise floor baseline (blue), which corresponds to the irreducible minimum MAE between noisy ground-truth values and a perfect reconstruction of the underlying noise-free signal (obtained using Eq. (22) with $\sigma=0$ and the ground-truth phase $\psi$).
IV.3 Assessing MPS generalization to out-of-distribution instances
The ability of a trained model to usefully extrapolate beyond its training distribution is captured by its out-of-distribution (OOD) performance. To determine the extent to which MPSTime generalizes usefully to OOD time series, we constructed a simple time-series imputation task that requires learning and extrapolating the dynamical rules of the observed data in order to demonstrate strong performance. We focused on the noisy trendy sinusoid (NTS), Eq. (22), with fixed trend ($m=1.0$), period ($\tau=30$), and Gaussian additive noise ($\sigma=0.1$). We constructed a restricted subset of training data ($1000$ NTS instances of length $T=100$ samples each), which encompass a small set of sinusoidal phases $\psi$ that can only take one of eight possible values ($\psi=\frac{\pi k}{4}$, where $k=0,1,\dots,7$ is an integer). MPSTime was trained on the restricted subset of data using a sufficiently large physical and maximum bond dimension, $d=10$ and $\chi_{\rm max}=60$, respectively. We then assessed the ability of MPSTime to accurately impute noisy sinusoids with phases that differ from those in the training set by uniformly discretizing the phase interval $[0,2\pi)$ into $5000$ points and generating an independent NTS test instance at each sampled phase. In each test instance, we masked a random contiguous block of 50% of the samples, and MPSTime was tasked with imputing the missing segment. We evaluated MPSTime against a noise-floor baseline, representing the irreducible error obtained when replacing a missing NTS segment with its noise-free counterpart (generated at the same test phase from the same NTS model in Eq. (22) with $\sigma=0$), and the same 1-NNI benchmark used above (in Sec. III.2).
Imputation performance, reported as mean absolute error (MAE), is shown as a function of phase for all test instances in Fig. 9. As expected, the noise floor baseline (shown in blue), yields a phase-independent MAE that captures the best achievable imputation performance in the absence of noise. For the 1-NNI benchmark, test instances are matched to their nearest neighbor, typically corresponding to a sinusoid in the training set with the closest phase $\psi$ to a given test instance. This behavior is reflected in the gold MAE$(\psi)$ curve in Fig. 9, which is minimal when the phase $\psi$ coincides with an example in the training set (vertical dashed lines), and increases with $\Delta\psi$ between a given test instance and the training set match. MPSTime OOD performance (red in Fig. 9) is qualitatively similar, with the MAE peaking at midpoints between the two nearest phases in the training set, achieves consistently lower MAE values than 1-NNI across the phase range. However, compared to 1-NNI, MPSTime has generally lower error, and matches the noise floor baseline at phases close to those contained in the training set (dotted vertical lines in Fig. 9), indicating that it has accurately imputed the underlying noise-free signal. Relative to 1-NNI, which always matches to a single noisy instance in the training set, this points to the ability of MPSTime to better encapsulate the deterministic component of the signal. This is likely due to our MPS imputation algorithm choosing the median of the RDM probability distribution (see Appendix B), which will tend to average over noise.
One way to interpret these findings is that MPSTime exhibits qualitatively similar behavior to 1-NNI through an MPS-based compression of the relevant temporal correlations contained in the training dataset. This interpretation is consistent with previous work by Martyn et al. [75] showing that instead of storing the entire training distribution of MNIST image data – a task requiring prohibitively large entanglement – an MPS learns a memory efficient low-entanglement representation, analogous to performing an efficient data compression. We note that although the MPS in this example cannot generalize information from training data of selected phases to learn all other phases, this property could easily be introduced into the MPS geometry using a translationally invariant MPS [76] which should, by construction, learn all phases from a single training phase.